OpenStudy (anonymous):

Solve each system of equations by substitution : y = –2x – 5 y = 3x + 20

OpenStudy (anonymous):

u need to get the value of x and y ?

OpenStudy (anonymous):

yes, please

OpenStudy (anonymous):

x=-5 and y=5

OpenStudy (anonymous):

Solve each system of equations by substitution. y = 2x – 11 y = –3x + 14

OpenStudy (anonymous):

3x + 20 +2x+5 = 0 5x = -25 x = -5 y = 3*(-5) +20 = 5

OpenStudy (anonymous):

x = 5 , y = -1

OpenStudy (anonymous):

from the first equation , substistute for y in the second. 3x+20 -2x-5 3x + 20 +2x+5 = 0 5x = -25 x = -5 y = 3*(-5) +20 = 5

OpenStudy (anonymous):

2 moar?

OpenStudy (anonymous):

oki

OpenStudy (anonymous):

its diff tho. 1 sec

OpenStudy (anonymous):

Simplify the rational expression: (x^2+4x-5)/(x^2-4x+3)

OpenStudy (anonymous):

factorise both the numerator and denominator...and cancel out the common factors.... u get (x+5)/(x-3)

OpenStudy (anonymous):

u sure the denominator u wrote is right ?

OpenStudy (anonymous):

oh nm

OpenStudy (anonymous):

1 mor

OpenStudy (anonymous):

Simplify the rational expression: (m+2)/(m^2-3m-10)

OpenStudy (anonymous):

1/(m-5)